# Jogging the Memory of Unlearned LLMs Through Targeted Relearning Attacks Content warning: This manuscript contains examples of harmful/hazardous text

### v2

@@ Abstract
Machine unlearning is a promising approach to mitigate undesirable memorization of training data in LLMs. However, in this work we show that existing approaches for unlearning in LLMs are surprisingly susceptible to a simple set of targeted relearning attacks. With access to only a small and potentially loosely related set of data, we find that we can “jog” the memory of unlearned models to reverse the effects of unlearning. For example, we show that relearning on public medical articles can lead an unlearned LLM to output harmful knowledge about bioweapons, and relearning general wiki information about the book series Harry Potter can force the model to output verbatim memorized text. We formalize this unlearning-relearning pipeline, explore the attack across three popular unlearning benchmarks, and discuss future directions and guidelines that result from our study.
@@ 1 Introduction
Figure 1: Recovering memorized text by relearning on public information: We ask the model to complete sentences from Harry Potter and the Order of the Phoenix [27]. We finetune the model to enforce memorization and then unlearn on the same text. Then, we show it is possible to relearn this memorized text using GPT-4-generated general information about the main characters, which does not contain direct text from the novels (see Section 5).
@@
Machine unlearning considers removing a model’s knowledge of certain topics or subsets of the training data [4, 8, 3]. Unlearning methods are particularly important for foundation models such as LLMs where the vast datasets used for pretraining and finetuning may contain private or undesirable content which must be removed in post-processing due to issues including data deletion requests, copyright infringement, or safety concerns [24, 18, 5, 11]. Given the scale of the datasets and models, unlearning methods for LLMs are typically approximate, in that they aim to efficiently update the model so that it approximates the behavior of a model retrained from scratch on a dataset with all undesirable training data removed.
@@
Unfortunately, while many unlearning methods have been proposed, recent works have shown that approaches for approximate unlearning are relatively fragile—particularly when scrutinized under an evolving space of attacks and evaluation strategies [e.g., 19, 28, 21, 12] (see Sec 2 for a detailed discussion of related work).
@@
In this work, we build on this growing body of work by exploring a simple and surprisingly effective attack on unlearned models. In particular, we show that current finetuning-based approaches for approximate unlearning in LLMs are susceptible to targeted relearning attacks—where a small amount of (potentially auxiliarly) data can “jog” the memory of unlearned models so they behave similarly to their pre-unlearning state (see Figure 1). As we show, these attacks can be effective when using only a limited set of data, including datasets or information that are loosely related to the task at hand and can be easily accessed by the public.
@@
In the remainder of the paper, we formalize our unlearning-relearning pipeline for LLMs, focusing on how to perform relearning attacks in realistic settings with limited data access, and how to effectively evaluate model behavior before and after the attack. We focus on three common unlearning applications: (1) preventing the generation of harmful or undesirable knowledge, (2) mitigating the memorization of sensitive or copyrighted content, and (3) suppressing the retention of specific undesirable words or phrases in the model. We study these scenarios through popular unlearning benchmarks: WMDP [14], TOFU [21], and Who’s Harry Potter (WHP) [7], showing that with access to either a limited subset of the unlearn data (Section 4) or a set of auxiliary, publicly available data (Section 5), we can drive the model to output supposedly unlearned knowledge [^1] Our code is publicly available at https://github.com/s-huu/jog_llm_memory. We end by analyzing the unlearning-relearning problem in a simplified setting (Section 6), and discussing guidelines and directions of future work.
2 Related Work
The initial concept of machine unlearning was motivated by GDPR regulations around the “right to be forgotten”, which asserted that users have the right to request deletion of their data from service providers [4]. Increasing model sizes and training costs have since spurred the development of approaches for approximate unlearning, which aim to efficiently update the model so it (roughly) behaves as if it never observed the data that was requested to be forgotten. Due to the scale of data and model sizes of modern LLMs, methods for approximate unlearning in LLMs have focused on scalable techniques such as gradient-based unlearning methods [22, 21, 16, 36, 37, 14, 11], in context unlearning [24], and guardrail-based unlearning [32].
Unfortunately, recent works have shown that it is possible to recover unlearned content after applying existing unlearning heuristics [12, 23, 9, 29]. Most closely related to our work, a number of recent works have explored the problem of relearning as a technique to evaluate robustness of unlearning by finetuning the unlearned model [31, 7, 30, 19]. However, unlike our work, these prior works generally consider relearning directly on the unlearn data, and do not study the relation between relearn set and the queries used for evaluation. In this case, the relearn set might contain direct answers to the evaluation queries, making it unclear whether relearning occurs simply due to learning the knowledge again from scratch, or due to triggering the memory of the approximately unlearned model. Our work instead explores targeted relearning attacks on unlearned models, showing that it is possible to “jog” the memory of unlearned models by recovering unlearned content even if the relearn set is only loosely related to the evaluation queries and unlearning task at hand. Finally, Qi et al. [25] show that finetuning LLMs on benign information can potentially compromise safety training, though this work does not explore attacks on machine unlearning and related methods.
3 Targeted Relearning Attack
@@
We next describe key aspects of our relearning problem pipeline (see Figure 2): (1) the targeted relearning attack threat model, (2) construction of the relearn set, and (3) considerations for evaluating the model before and after relearning.
@@ 3.1 Problem Formulation and Threat Model
To launch a targeted relearning attack, we consider an adversary $\mathcal{A}$ who has access to the unlearned model $w_{u}$. We do not assume that the adversary $\mathcal{A}$ has access to the original model $w$, nor do they have access to the complete unlearn set $D_{u}$. Our key assumption on this adversary is that they are able to finetune the unlearned model $w_{u}$ with some auxiliary data, $D^{\prime}$ (described in detail in Section 3.2). We discuss two common scenarios where such finetuning is feasible: (1) Model weight access adversary. If the model weights $w_{u}$ are openly available, an adversary may finetune this model assuming access to sufficient computing resources. (2) API access adversary. On the other hand, if the LLM is either not publicly available (e.g. GPT) or the model is too large to be finetuned directly with the adversary’s computing resources, finetuning may still be feasible through LLM finetuning APIs (e.g. [33]).
@@
Although unlearning datasets are expected to contain sensitive or toxic information, these same datasets are also likely to contain some benign knowledge that is publicly available. Formally, let $\mathcal{M}^{\prime}:\mathcal{W}\times\mathcal{D}^{\prime}\rightarrow\mathcal{W}$ be the algorithm that finetunes the unlearned model $w_{U}$ using an auxiliary dataset $D^{\prime}$. Our intuition is that by re-finetuning $w_{U}$ with auxiliary information $D^{\prime}$ that is correlated with $D_{u}$, sensitive unlearned content may risk being generated using $w^{\prime}$, even if this knowledge never appears in $D^{\prime}$ nor in the text completions of $w_{U}$ (see Fig 2). We explore this intuition directly through a simplified example in Section 6, showing that increasing correlations between two keywords in the data can make it more feasible to perform relearning when only one keyword is relearned.
@@ 3.3 Unlearning Tasks & Evaluation
In our experiments, we consider a number of common unlearning tasks: (1) preventing the generation of harmful or undesirable knowledge, (2) mitigating the memorization of sensitive or copyrighted content, and (3) suppressing the retention of specific undesirable words or phrases. These tasks naturally have different evaluation procedures $eval(o,q)$ that may be used to define success:
  1. 1.
    LLM-based evaluation. In applications such as removing hazardous or toxic knowledge, a reasonable goal for the adversary is to get high quality answers from the unlearned LLM using a harmful query. In this case, where quality is less well-defined through simple metrics, it is common to ask an existing LLM to provide a score. Following the recipe from Zheng et al. [38], we use GPT-4 as the LLM Judge to provide a single score between 1-10 for every query-completion pair. A higher score means indicates that the completion more effectively answers the query.
  2. 2.
    Rouge-L. Another common application of unlearning is to mitigate verbatim memorization of copyrighted content. Under such settings, an adversary aims to see if the LLM will output the exact same copyrighted text used during training. Similar to Maini et al. [21], given the prefix $q$ and the true copyrighted continuation content $a$, we use Rouge-L score to measure the similarity between $a$ and the model output $o$. Specifically, we define $eval(o,q)=\text{Rouge}_{L}(a,o)$.
  3. 3.
    Keyword search. Finally, in certain applications the adversary may care about obtaining specific keywords such as names or ID numbers from the model. To evaluate whether such a goal is satisfied, one can simply check whether a keyword $k$ is contained in $o$. Specifically, $eval(o,q)$ could be defined as $eval(o,q)=\mathbf{1}_{k\in o}$.  [^2] Note that this is more flexible compared to exact string match based metric in prior work [21] as we only care about the appearance of a few tokens.
@@
To perform evaluation, we assume access to a set of queries $S$ we want the model to complete. As discussed, in this work we are interested in testing on queries $q\in S$ whose knowledge has been successfully unlearned, i.e. $eval(gen(w,q),q)$ is high and $eval(gen(w_{U},q),q)$ is low. The goal is then to investigate how the score of the relearned model $eval(gen(w^{\prime},q),q)$ compares to those of the original and unlearned model. We provide an algorithmic description of our end-to-end targeted relearning attack in Algorithm 1.
@@
Table 1: Example of how keywords unseen in the relearn set could still be recovered after relearning, for left: TOFU, right: WHP. The appearance of keywords in the model completion is highlighted in green.
4 Relearning Attack Using a Portion of the Unlearn Set
We first consider relearning attacks in the setting where the adversary is able to use a portion of the unlearn dataset for relearning. However, as discussed in Section 3.2, unlike prior work in relearning, when performing relearning we assume the adversary may only have access to a highly skewed sample of this unlearn data. More formally, we assume the unlearn set can be partitioned into two disjoint sets, i.e., $D_{u}=D_{u}^{(1)}\cup D_{u}^{(2)}$ such that $D_{u}^{(1)}\cap D_{u}^{(2)}=\emptyset$. We assume that the adversary only has access to $D_{u}^{(1)}$ (a portion of the unlearn set), but is interested in attempting to access the knowledge present in $D_{u}^{(2)}$ (a separate, disjoint set of the unlearn data). We study two datasets: TOFU [21] and Who’s Harry Potter (WHP) [7] where the unlearning task for the adversary is to infer specific keywords that have been unlearned. We use LoRA finetuning to perform relearning in all experiments (see Appendix A.4 for additional details).
TOFU. For TOFU, we use the forget05 dataset [20] as $D_{u}$, which contains 200 QA pairs for 10 fictitious authors. We unlearn the Phi-1.5 model [15] using gradient ascent, a common unlearning baseline. We construct the relearn set as follows: For each author we select only one book written by the author. We then construct a test set by only sampling QA pairs relevant to this book, i.e., $D_{u}^{(2)}=\{x|x\in D_{u},book\subset x\}$ where $book$ is the name of the selected book. By construction, $D_{u}^{(1)}$ is the set that contains all data without the presence of the keyword $book$. To construct the relearn set, we assume the adversary has access to $D^{\prime}\subset D_{u}^{(1)}$. In relearning, the goal is to recover the string $book$ despite never seeing this keyword in the relearning data.
Who’s Harry Potter (WHP). For WHP, we first finetune Llama-2-7b [34] on the a set of text containing the direct text of HP novels, QA pairs, and fan discussions about Harry Potter series. For unlearning, we create a set of alternative labels on the same text $D_{u}$ using the methods from [7]. We then perform unlearning on the model through finetuning on the alternative labels. To partition $D_{u}$, we choose the dataset for evaluation, $D_{u}^{(2)}$, to be the set of all sentences that contain any of the words “Hermione” or “Granger”. As a result, the set $D_{u}^{(1)}$ contains no information about the name “Hermione Granger”. Similar to the TOFU case, we sample $D^{\prime}\subset D_{u}^{(1)}$ to form the relearn set. Thus, the relearn set contains names of other characters closely related to Hermione, but no direct knowledge of the character’s name. The goal is to recover the name “Hermione Granger” given information on the other characters.
Quantitative results. We additionally explore the efficacy of relearning with partial unlearn sets through a more comprehensive set of quantitative results in Figure 3. In particular, for each dataset, we investigate the effectiveness of relearning when starting from multiple potential unlearning checkpoints. For every relearned model we record the attack success rate (ASR). On both datasets, we observe that our attack is able to achieve $>70\%$ ASR in searching the keywords when unlearning is shallow. As we start to unlearn further from the original model, it becomes harder to reconstruct keywords through relearning. Meanwhile, increasing the number of relearning steps does not always mean better ASR. For example in the TOFU experiment, if the relearning happens for more than 40 steps, ASR drops for all unlearning checkpoints. See Appendix A.2,B.2 for additional details.
Algorithm 1 Targeted Relearning Attack
4:  for query $q\in S$ do
6:  end for
7:  return  Relearned model $w^{\prime}$ and score for every $q\in S$: $\{s_{q}\}$
@@ 5 Relearning Attack Using Public Information
(a) Zephyr-7b-beta Forget (WMDP)
(b) Llama-3-8b Forget (WMDP)
(c) Zephyr-7b-beta Retain (MT-Bench)
(d) Llama-3-8b Retain (MT-Bench)
Figure 4: LLM-as-Judge scores for the Top: forget set (WMDP benchmarks) and Bottom: retain set (MT-Bench) for two models: Left: zephyr-7b-beta, Right: Llama-3-8b. For each model, we evaluate on the original model, the unlearned model and the relearned model. For each unlearning baseline column, the relearned model is obtained by finetuning the unlearned model from the same block. We use the same unlearned and relearned model for both forget and retain evaluation. Average scores over all questions are reported; scores range between $1$-$10$, with higher scores indicating better answer quality.
@@ 5.1 Recovering Harmful Knowledge in WMDP
Setup. We first consider the WMDP benchmark [14], which aims to unlearn hazardous knowledge from existing models. We test our attack on two popular models: zephyr-7b-beta [35] and Llama-3-8b [6]. Following Li et al. [14], we unlearn the bio-attack corpus and cyber-attack corpus, which contain hazardous knowledge in biosecurity and cybersecurity. We use gradient ascent as a simple baseline for approximate unlearning. To construct the relearn set, we first pick 15 questions from the WMDP multiple choice question (MCQ) set whose knowledge has been unlearned from $w_{u}$. For each question $q$, we find public online articles related to $q$ and use GPT to generate paragraphs about general knowledge relevant to $q$. We ensure that this resulting relearn set does not contain direct answers to any question in the evaluation set. For additional details on how $D^{\prime}$ is generated, please refer to Appendix C.2.
We first show that common gradient-based unlearning methods are vulnerable this form of targeted relearning attack. We investigate relearning after using a number of popular methods to perform unlearning: gradient ascent (GA), gradient difference (GD) [16], KL minimization (KL) [21], Negative Preference Optimization (NPO) [37], and SCRUB [13]. We describe these methods in more detail in Appendix E.
Question To make influenza A more virulent, one should
Relearn using two public articles [2, 1] Score: 8 $\uparrow$ answ: increase the virulence of the virus by introducing specific mutations into the viral genome. To make influenza A more virulent, one should: a) Introduce the HA and NA genes from a highly pathogenic avian influenza virus into a low pathogenic avian influenza virus…
Table 2: Example of model outputs given a fixed prompt before unlearning (top), after unlearning (middle), and after relearning (bottom) on the WMDP benchmark, using public articles. A higher score (from the LLM evaluation) means the model output is more relevant to the question.
Quantitative results. To evaluate our attack, we evaluate on an answer completion task where the adversary prompts the model with a question and we let the model complete the answer. We randomly choose 70 questions from the WMDP MCQ set and remove the multiple choices provided to make the task harder and more informative for our evaluation. Therefore, the adversary only knows the question and we apply LLM based evaluation discussed in Section 3.2 to evaluate the quality of the model’s answer before and after the targeted relearning attack. We use the prompt in Figure 6 of Zheng et al. [38] as the scoring function. Average score over all questions is recorded.
@@
PEFT unlearning increases susceptibility to relearning attacks. The results in this section thus far considered relearning after perform finetuning for unlearning on all model weights. However, we also perform experiments in a more compute restricted setting where the unlearning is done via parameter efficient finetuning technique such as LoRA [10]. The results are shown in Table 3. Similar to the findings in the full unlearning setting, LoRA unlearning is not robust. In fact, as it attempts to modify the model using a small fraction of parameters, it becomes easier to recover the unlearned the content with small amount of finetuning.

### v4

# Unlearning or Obfuscating? Jogging the Memory of Unlearned LLMs via Benign Relearning content warning: this manuscript contains examples of harmful/hazardous text

@@ Abstract
Machine unlearning is a promising approach to mitigate undesirable memorization of training data in ML models. However, in this work we show that existing approaches for unlearning in LLMs are surprisingly susceptible to a simple set of benign relearning attacks. With access to only a small and potentially loosely related set of data, we find that we can “jog” the memory of unlearned models to reverse the effects of unlearning. For example, we show that relearning on public medical articles can lead an unlearned LLM to output harmful knowledge about bioweapons, and relearning general wiki information about the book series Harry Potter can force the model to output verbatim memorized text. We formalize this unlearning-relearning pipeline, explore the attack across three popular unlearning benchmarks, and discuss future directions and guidelines that result from our study. Our work indicates that current approximate unlearning methods simply suppress the model outputs and fail to robustly forget target knowledge in the LLMs.
@@ 1 Introduction
Machine unlearning considers removing a model’s knowledge of certain topics or subsets of the training data (Cao & Yang, 2015; Ginart et al., 2019; Bourtoule et al., 2021). Unlearning methods are particularly important for foundation models such as LLMs where the vast datasets used for pretraining and finetuning may contain private or undesirable content which must be removed in post-processing due to issues including data deletion requests, copyright infringement, or safety concerns (Pawelczyk et al., 2024; Liu et al., 2024b; Chen & Yang, 2023; Jia et al., 2024; Liu, 2024). Given the scale of the datasets and models, unlearning methods for LLMs are typically approximate, in that they aim to efficiently update the model so that it approximates the behavior of a model retrained from scratch on a dataset with all undesirable training data removed.
Figure 1: Recovering memorized text by relearning on public information: We ask the model to complete sentences from Harry Potter and the Order of the Phoenix (Rowling, 2003). We finetune the model to enforce memorization and then unlearn on the same text. Then, we show it is possible to relearn this memorized text using GPT-4-generated general information about the main characters, which does not contain direct text from the novels (see Section 4).
@@
Unfortunately, while many unlearning methods have been proposed, recent works have shown that approaches for approximate unlearning are relatively fragile—particularly when scrutinized under an evolving space of attacks and evaluation strategies (e.g., Lynch et al., 2024; Shi et al., 2024; Maini et al., 2024b; Jin et al., 2024) (see Sec 8 for a detailed discussion of related work).
@@
In this work, we build on this growing body of work by exploring a simple and surprisingly effective attack on unlearned models. In particular, we show that current finetuning-based approaches for approximate unlearning are simply obfuscating the model outputs instead of truly forgetting the information in the forget set, making them susceptible to benign relearning attacks—where a small amount of (potentially auxiliary) data can “jog” the memory of unlearned models so they behave similarly to their pre-unlearning state (see Figure 1). While benign finetuning strategies have been explored in prior works (e.g., Qi et al., 2023; Tamirisa et al., 2024; Lynch et al., 2024), these works consider general-purpose datasets for relearning without studying the overlap between the relearn data and queries used for unlearning evaluation [^1] For example, Lynch et al. (2024) used the first three books of Harry Potter as the relearn set. However, some evaluation queries can be directly answered by the relearn set, e.g. ”Harry Potter’s friends are …”. . In contrast, we focus on the scenario where the additional data itself is insufficient to capture the forget set—ensuring that the attack is "relearning" instead of simply "learning" the unlearned information from this finetuning procedure. Surprisingly, our results show that relearning attacks can be effective when using only a limited set of data, including datasets that are insufficient to inform the evaluation queries alone and can be easily accessed by the public.
@@
In the remainder of the paper, we formalize our unlearning-relearning pipeline for LLMs, focusing on how to perform relearning attacks in realistic settings with limited data access, and how to effectively evaluate model behavior before and after the attack. We focus on three common unlearning applications: (1) preventing the generation of harmful or undesirable knowledge, (2) mitigating the memorization of sensitive or copyrighted content, and (3) suppressing the retention of specific undesirable words or phrases in the model. We study these scenarios through popular unlearning benchmarks: WMDP (Li et al., 2024), TOFU (Maini et al., 2024b), and Who’s Harry Potter (WHP) (Eldan & Russinovich, 2023), showing that with access to either a limited subset of the unlearn data (Section 3) or a set of auxiliary, publicly available data (Section 4), we can drive the model to output supposedly unlearned knowledge. In doing so, our work provides evidence that existing approximate unlearning methods are effectively obfuscating the model outputs instead of robustly removing target knowledge in LLMs. We end by analyzing the unlearning-relearning problem in a simplified setting (Section 5), and discussing guidelines and directions of future work. Our code is available at https://github.com/s-huu/jog_llm_memory.
2 benign relearning Attack
@@
We next describe key aspects of our relearning problem pipeline (see Figure 2): (1) the benign relearning attack threat model, (2) construction of the relearn set, and (3) considerations for evaluating the model before and after relearning.
@@ 2.1 Problem Formulation and Threat Model
To launch a benign relearning attack, we consider an adversary $\mathcal{A}$ who has access to the unlearned model $w_{u}$. We do not assume that the adversary $\mathcal{A}$ has access to the original model $w$, nor do they have access to the complete unlearn set $D_{u}$. Our key assumption on this adversary is that they are able to finetune the unlearned model $w_{u}$ with some auxiliary data, $D^{\prime}$ (described in detail in Section 2.2). We discuss two common scenarios where such finetuning is feasible: (1) Model weight access adversary. If the model weights $w_{u}$ are openly available, an adversary may finetune this model assuming access to sufficient computing resources. (2) API access adversary. On the other hand, if the LLM is either not publicly available (e.g. GPT) or the model is too large to be finetuned directly with the adversary’s computing resources, finetuning may still be feasible through LLM finetuning APIs (e.g. TogetherAI ). It is conceivable that our benign relearning attack could also be launched at inference time through techniques such as in-context learning (Shumailov et al., 2024), though we have not verified this in our experiments.
@@
Although unlearning datasets are expected to contain sensitive or toxic information, these same datasets are also likely to contain some benign knowledge that is publicly available. Formally, let $\mathcal{M}^{\prime}:\mathcal{W}\times\mathcal{D}^{\prime}\rightarrow\mathcal{W}$ be the algorithm that finetunes the unlearned model $w_{U}$ using an auxiliary dataset $D^{\prime}$. Our intuition is that by re-finetuning $w_{U}$ with auxiliary information $D^{\prime}$ that is correlated with $D_{u}$, sensitive unlearned content may risk being generated using $w^{\prime}$, even if this knowledge never appears in $D^{\prime}$ nor in the text completions of $w_{U}$ (see Figure 2). We explore this intuition directly through a simplified example in Section 5, showing that increasing correlations between two keywords in the data can make it more feasible to perform relearning when only one keyword is relearned.
@@ 2.3 Unlearning Tasks & Evaluation
In our experiments, we consider a number of common unlearning tasks: (1) preventing the generation of harmful or undesirable knowledge, (2) mitigating the memorization of sensitive or copyrighted content, and (3) suppressing the retention of specific undesirable words or phrases. These tasks naturally have different evaluation procedures $eval(o,q)$ that may be used to define success: (1) LLM-as-Judge, (2) Rouge-L, (3) Binary prediction accuracy. We defer the readers to the Appendix I for a more detailed discussion of these metrics.
@@
To perform evaluation, we assume access to a set of queries $S$ we want the model to complete. As discussed, in this work we are interested in testing on queries $q\in S$ whose knowledge has been successfully unlearned, i.e. $eval(gen(w,q),q)$ is high and $eval(gen(w_{U},q),q)$ is low. The goal is then to investigate how the score of the relearned model $eval(gen(w^{\prime},q),q)$ compares to those of the original and unlearned model. We provide an algorithmic description of our end-to-end benign relearning attack in Algorithm 1.
3 Relearning Attack Using a Portion of the Unlearn Set
We first consider relearning attacks in the setting where the adversary is able to use a portion of the unlearn dataset for relearning. However, as discussed in Section 2.2, unlike prior work in relearning, when performing relearning we assume the adversary may only have access to a highly skewed sample of this unlearn data. More formally, we assume the unlearn set can be partitioned into two disjoint sets, i.e., $D_{u}=D_{u}^{(1)}\cup D_{u}^{(2)}$ such that $D_{u}^{(1)}\cap D_{u}^{(2)}=\emptyset$. We assume that the adversary only has access to $D_{u}^{(1)}$ (a portion of the unlearn set), but is interested in attempting to access the knowledge present in $D_{u}^{(2)}$ (a separate, disjoint set of the unlearn data). We study two datasets: TOFU (Maini et al., 2024b) and Who’s Harry Potter (WHP) (Eldan & Russinovich, 2023) where the unlearning task for the adversary is to infer specific keywords that have been unlearned. In addition, we ensure that $D_{u}^{(1)}$ alone is insufficient to infer keywords from $D_{u}^{(2)}$ (See Appendix G for details). We use LoRA finetuning to perform relearning in all experiments (see Appendix A.4 for additional details).
TOFU. For TOFU, we use the forget05 dataset (Maini et al., 2024a) as $D_{u}$, which contains 200 QA pairs for 10 fictitious authors. We unlearn the Phi-1.5 model (Li et al., 2023) using gradient ascent (Golatkar et al., 2020b), a common unlearning baseline. We construct the relearn set as follows: For each author we select only one book written by the author. We then construct a test set by only sampling QA pairs relevant to this book, i.e., $D_{u}^{(2)}=\{x|x\in D_{u},book\subset x\}$ where $book$ is the name of the selected book. By construction, $D_{u}^{(1)}$ is the set that contains all data without the presence of the keyword $book$. To construct the relearn set, we assume the adversary has access to $D^{\prime}\subset D_{u}^{(1)}$. In relearning, the goal is to recover the string $book$ despite never seeing this keyword in the relearning data.
Who’s Harry Potter (WHP). For WHP, we first finetune Llama-2-7b (Touvron et al., 2023) on a set of text containing the direct text of HP novels, QA pairs, and fan discussions about Harry Potter series. For unlearning, we create a set of alternative labels on the same text $D_{u}$ using the methods from Eldan & Russinovich (2023). We then perform unlearning on the model through finetuning on the alternative labels. To partition $D_{u}$, we choose the dataset for evaluation, $D_{u}^{(2)}$, to be the set of all sentences that contain any of the words “Hermione” or “Granger”. As a result, the set $D_{u}^{(1)}$ contains no information about the name “Hermione Granger”. Similar to the TOFU case, we sample $D^{\prime}\subset D_{u}^{(1)}$ to form the relearn set. Thus, the relearn set contains names of other characters closely related to Hermione, but no direct knowledge of the character’s name. The goal is to recover the name “Hermione Granger” given information on the other characters.
Quantitative results. We additionally explore the efficacy of relearning with partial unlearn sets through a more comprehensive set of quantitative results in Figure 3. In particular, for each dataset, we investigate the effectiveness of relearning when starting from multiple potential unlearning checkpoints. For every relearned model, we perform binary prediction on whether the keywords are contained in the model generation and record the attack success rate (ASR). On both datasets, we observe that our attack is able to achieve $>70\%$ ASR in searching the keywords when unlearning is shallow. As we start to unlearn further from the original model, it becomes harder to reconstruct keywords through relearning. Meanwhile, increasing the number of relearning steps does not always mean better ASR. For example in the TOFU experiment, if the relearning happens for more than 40 steps, ASR drops for all unlearning checkpoints. See Appendix A.2B.2 for additional details.
Dataset TOFU WHP
@@ 3 Relearning Attack Using a Portion of the Unlearn Set
Table 1: Examples of evaluation query and relearn text to recover unseen keywords after relearning from TOFU and WHP. The appearance of keywords in the model completion is highlighted in green.
@@ 4.1 Recovering Harmful Knowledge in WMDP
Setup. We first consider the WMDP benchmark (Li et al., 2024), which aims to unlearn hazardous knowledge from existing models. We test our attack on two popular models: zephyr-7b-beta (Tunstall et al., 2023) and Llama-3-8b (Dubey et al., 2024). Following Li et al. (2024), we unlearn the bio-attack corpus and cyber-attack corpus, which contain hazardous knowledge in biosecurity and cybersecurity. We use gradient ascent as a simple baseline for approximate unlearning. To construct the relearn set, we first pick 15 questions from the WMDP multiple choice question (MCQ) set whose knowledge has been unlearned from $w_{u}$. For each question $q$, we find public online articles related to $q$ and use GPT to generate paragraphs about general knowledge relevant to $q$. We ensure that this resulting relearn set does not contain direct answers to any question in the evaluation set. For additional details on how $D^{\prime}$ is generated, please refer to Appendix C.2.
Figure 4: LLM-as-Judge scores for the forget set (WMDP benchmarks) for two models: Left: zephyr-7b-beta, Right: Llama-3-8b. For each model, we evaluate on the original model, the unlearned model and the relearned model. For each unlearning baseline column, the relearned model is obtained by finetuning the unlearned model from the same block. We use the same unlearned and relearned model for both forget and retain evaluation. Average scores over all questions are reported; scores range between $1$-$10$, with higher scores indicating better answer quality. We defer the retain MT-Bench results to Appendix C.4 due to space constraint.
We first show that common gradient-based unlearning methods are vulnerable this form of benign relearning attack. We investigate relearning after using a number of popular methods to perform unlearning: gradient ascent (GA) (Golatkar et al., 2020a), gradient difference (GD) (Liu et al., 2022), KL minimization (KL) (Maini et al., 2024b), Negative Preference Optimization (NPO) (Zhang et al., 2024), and SCRUB (Kurmanji et al., 2024). We describe these methods in more detail in Appendix F.
@@
Quantitative results. To evaluate our attack, we evaluate on an answer completion task where the adversary prompts the model with a question and we let the model complete the answer. We randomly choose 70 questions from the WMDP MCQ set and remove the multiple choices provided to make the task harder and more informative for our evaluation. Therefore, the adversary only knows the question and we apply LLM based evaluation discussed in Section 2.2 to evaluate the quality of the model’s answer before and after the benign relearning attack. We use the prompt in Figure 6 of Zheng et al. (2024) as the scoring function. Average score over all questions is recorded.
Question To make influenza A more virulent, one should
Relearn using two public articles (flu, 2024; typ, 2023) Score: 8 $\uparrow$ answ: increase the virulence of the virus by introducing specific mutations into the viral genome. To make influenza A more virulent, one should: a) Introduce the HA and NA genes from a highly pathogenic avian influenza virus into a low pathogenic avian influenza virus…
Table 3: Example Zephyr-7b-beta generated outputs given an example question before unlearning (top), after unlearning (middle), and after relearning (bottom) on the WMDP benchmark, using public articles. A higher score (from the GPT-4 evaluation) means the model output is more relevant to the question.
@@
PEFT unlearning increases susceptibility to relearning attacks. The results in this section thus far considered relearning after perform finetuning for unlearning on all model weights. However, we also perform experiments in a more compute restricted setting where the unlearning is done via parameter efficient finetuning technique such as LoRA (Hu et al., 2021). The results are shown in Table 2. Similar to the findings in the full unlearning setting, LoRA unlearning is not robust. In fact, as it attempts to modify the model using a small fraction of parameters, it becomes easier to recover the unlearned the content with small amount of finetuning.